UT 20160425 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson,
Observer AO: DMiller, GTaylor (both LBTO), EPinna, APuglisi (remote from Italy upon request)
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

- Stability FLAO+LUCI1, AFC, Stability PSF

Summary

- AFC seems to work in broad-band filters

- Bino-script for photometric stds LUCi1/2 successfully executed

- 2/3 night lost due to wind and humidity

- AFC-tests closed dome

Details

All times are UT
Files are luci1.20160425.NNNN.fits

01:50 Steve's opening up

02:00 Sunset. Weather: photometric, T = +4deg, Hum = 35%, wind ~ 10-15 m/s (need to stay away from SW)

02:20 Steve checks pointing and collimates, DIMM reports 1.5"-2" frown, sad smile

Photometric std LUCI1+2

02:39 Preset to S9142_phot_bin, OK!, N30, KKsHJz, #0013-0022 (L1), #0003-0012 (L2), DIMM ~ 1.0-1.5"
N3.75, zJHKsK, #0023-0032 (L1), #0013-0022 (L2), seing 0.7-1.0" on N3.75-camera images
Time to completion: 42min for 10*1min + 10*10sec (12min) for both cameras incl. mask exchange = 29% Duty Cycle
SUPER, full SL Bino imaging-script made it without problems

AFC

03:23 Preset to SDSS1016, OK, RunAO failed, didn't find the RS (but its definitively a star) but rather tried to lock on something else, preset again,
RunAO, OK, RS has 15.1,Bin=4, 36 modes, 200Hzm #0033 AFC-acquisition ima, #0034 verification, need to be optimized,
#0035 OK, Ks, #0036 (5min), AFC not perfect, #0037 AFC-acq, #0038-0039 AFC-verification, #0040-0045 (5min), DIMM 1.0"
AO-worked, AFC seems to work in BB as well

04:22 Preset to NGC4147_AFC, OK, RunAO, OK, RS=12.6, Bin=3, 66mode, 200Hz, ACQ-acq,K, #0046,
skipped, H: AFC-acq #0047, ACQ-veri #0048, AFC #0049-0051 (5min), #0049 vs #0050 not very good,
#0050 vs #0051 OK, J: AFC-acq #0052, AFC #0053-0055, need to face off the wind as we are very close to the limit

05:20 Need to close due to wind (again)

Alexander does closed-dome tests

09:30 Still closed due to wind (and humidity in the meantime)

12:00 End of night, still closed for wind and humidity

-- %USERSIG{JochenHeidt - 2016-04-25}%
Topic revision: r5 - 25 Apr 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback